Calculate the 95% prediction interval of y when x=5 using the 2000 pairs

Mean of x = 4.51

Regression Statistics
Multiple R 0.012848
R Square 0.000165
Adjusted R Square -0.00034
Standard Error 2.869737
Observations 2000
ANOVA
df SS MS F Significance F
Regression 1 2.716416 2.716416 0.329847 0.565814
Residual 1998 16454.31 8.235388
Total 1999 16457.02
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 4.509054 0.119572 37.70997 1.7E-235 4.274555 4.743552574 4.274555 4.743553
X 0.012884 0.022434 0.574323 0.565814 -0.03111 0.056879632 -0.03111 0.05688
Answer #1

The regression model is,

\hat{y} = 4.509054 + 0.012884 X

For x = 5,

\hat{y} = 4.509054 + 0.012884 * 5 = 4.573474

Standard error of regression, s = 2.869737

DF Residual = 1998

As, DF Residual is large, critical t value for 95% prediction interval is 1.96

SS Regression = SSxy^2 / SSxx = 2.716416

Slope = SSxy / SSxx = 0.012884

SSxy = 2.716416 / 0.012884 = 210.8364

SSxx = 210.8364 / 0.012884 = 16364.2

95% prediction interval of y when x=5 is,

\hat{y}\pm t * S * \sqrt{1 + (1/n) + (x - \bar{x})^2 / SS_{xx}}

4.573474 \pm 1.96 * 2.869737 * \sqrt{1 + (1/2000) + (5 - 4.51)^2 / 16364.2}

= 4.573474  \pm  5.62613

= (4.573474 - 5.62613 ,  4.573474 + 5.62613)

= (-1.052656 , 10.1996)